your younger brother is getting stickers from a vending machine at the local supermarket. a proportion of stickers have the phrase "girls rule." if there are 8 stickers with this message in a sample of 36 stickers, what is the 95% confidence interval for this proportion


1
Expert's answer
2021-11-05T15:06:43-0400

"\hat p=\frac{8}{36}=0.2222."

"95\%CI=(0.2222-1.96\sqrt{\frac{0.2222(1-0.2222)}{36}},0.2222+1.96\sqrt{\frac{0.2222(1-0.2222)}{36}})="

"=(0.0864,0.3580)."